Case Study 2-2: Lacto-ovovegetarian Diet Planning Sarah T. is a 20-year-old college student who does not eat meat, fish, or poultry. She is 5 feet 7 inches tall, weighs 140 pounds, and is physically active most days, riding her bike to school from her apartment off campus. Although her weight has been stable for the past year, Sarah’s mother is concerned she is not getting the nutrients she needs for optimal health. Sarah’s usual daily diet includes a toaster pastry and juice for breakfast, peanut butter sandwich for lunch, and pasta or vegetable pizza for dinner. She snacks frequently on chips or cookies and drinks one or two diet sodas each day. 1. Using the glossary in Highlight 2 of this chapter, how would you categorize Sarah’s diet? 2. What key nutrients are likely to be inadequate in Sarah’s current diet? 3. What changes or additions to her diet would you recommend to include sources of these nutrients? 4. What is a reasonable estimate of Sarah’s daily kcalorie needs? Based on this estimate, what is the daily amount of protein foods (in ounces) that you would recommend for Sarah? 5. Using Table H2-1 as your guide, set up a 7-day plan to show how Sarah can meet her recommended daily protein needs.
Answer #1

1. According to the case study diet of Sarah is Lacto -ovovegetarian. Lacto -ovo vegetarians are like vegetarians who consumes green vegetables, whole grains, fruits and they also consume milk products and eggs .

2. Sarah according to her diet she donot eat meat, fish or poultry but , she consumes eggs & milk products along with vegetables and fruits. She may be inadequate is protein, zinc, iron, calcium, vitamin B12, vitamin D and omega 3 fatty acids.

3.Sarah has to be adviced to consume adequate amount of vegetables which provide iron and calcium, whole-grains and cereals which fulfill her protein, iron, zinc requirements.She can include dried fruit and nuts or seeds as snacks which will satisfy her iron, calcium requirements. Nuts and soybeans will be a good supplement of omega 3 fatty acids which is very important for cardiovascular health. Adequate eggs consume can give her the necessary protein needed.

4. BMI of Sarah .Formula =703 x weight (lbs) / [height (in)]2   

=703 x 140/ (67)2

=21.9

5feet = 60 inches

5feet 7 inches= 67 inches.

Weight in kg= 140/2.2= 63kgs

BMR(basal metabolic rate)= 4.7 x (height in inches )+(4.35 x weight in pounds )- (4.7 x age in years ) +655

= (4.7x 67)+ (4.35x140)-(4.7x20+655)= 314+609-94= 829 +655=1484(BMR)

Lean mass percentage for an active female = 1.65

Sarah's estimated kcal need = 1484 x 1.65=2448 kcal

Her estimated daily kcal needs 2448. Recommended daily amounts for 2448 kcal.

Her need for carbohydrates = 315 to 378 grams /day (Moderate intensity training person carbohydrate need is 5-6 grams/kg)

Her need for proteins =100 to 126 grams/day( Protein need for active person =1.6 - 2 grams of protein /kg)

Need of protein is about 4.5 ounces per day for Sarah.